# Matrix harmonic analysis at high temperature via the Dirichlet process

### v1

@@ 3.4. Weak convergence in the high temperature regime
Substituting the two inequalities in Lemma 3.9 and Lemma 3.10 into (3.55) gives
$\begin{split}\left|F_{N}(t)-F(t)\right|&\leq\int_{\mathbb{R}}\mathrm{d}x(C_{T,\delta})^{cM_{T,\rho}}(x^{2}+\delta^{2})^{-cM_{T,\rho}/2+|c_{N}-c|/2}\left(b_{N}+\frac{|c_{N}-c|}{2}\log(x^{2}+\delta^{2})\right)e^{b_{N}}\end{split}$ (3.73)
For large enough $N$, the difference $|c_{N}-c|<\varepsilon$ can be bounded by any small $\varepsilon$. Thus
$cM_{T,\rho}/2+|c_{N}-c|/2<-1/2$ (3.74)
can be achieved, and so the previous integral is absolute integrable. Noting that $b_{N}\to 0$ and $|c_{N}-c|\to 0$ as $N\to\infty$ results $|F_{N}(t)-F(t)|\to 0$, which finishes the proof. ∎
3.5. Proofs for Theorem 1.1 and 1.2
@@
and $\rho_{N}^{(c_{N})}$ is the distribution of the random mean of the Dirichlet process with parameter measure $c_{N}\rho$. By Section 3.3, in the limit where $c_{N}:=N\beta_{N}/2\to 0$ as $N\to\infty$, the measure $\rho_{N}^{(c_{N})}$ converges weakly to $\rho$. Thus, by the dominated convergence theorem,
$\lim_{N\to\infty}\int_{\mathbb{R}}e^{ux}\rho_{N}^{(c_{N})}(\mathrm{d}x)=\int_{\mathbb{R}}e^{ux}\rho(\mathrm{d}x),$ (3.77)
@@
Now for (1.21), its proof is almost identical, except that in the $N\to\infty$ one has instead $c_{N}:=N\beta_{N}/2\to c>0$. Then by Section 3.4, one has the weak convergence $\rho_{N}^{(c_{N})}\to\rho^{(c)}$. Then applying the dominated convergence theorem gives us the result.
@@ 4.1. Comparison with results in [3]
As mentioned in the introduction section, Benaych-Georges, Cuenca and Gorin [3] develop the notion of Bessel generating function of a symmetric probability measure $\mu$ on $\mathbb{R}^{N}$, which is an integral transform with the multivariate Bessel function as the kernel:
$G(\vec{x};\mu):=\int_{\mathbb{R}^{N}}\mathcal{B}_{\vec{a}}(\vec{u};\theta)\mu(\mathrm{d}\vec{a}).$ (4.1)
@@
Here we will restate their results in the following special case: when the Bessel generating function has only one non-zero variable and $\mu_{N}$ is (symmetrically) finitely supported on $N$ points $a_{1},\ldots,a_{N}$, we have
$G(\underbrace{x,0,\ldots,0}_{N\text{ entries}};\mu)=\int_{\mathbb{R}^{N}}\mathcal{B}_{\vec{b}}(x;N,\theta)\prod_{j=1}^{N}\delta_{a_{j}}(\mathrm{d}b_{j})=\mathcal{B}_{\vec{a}}(x;N,\theta).$ (4.2)
Proposition 4.1 (Special case of [3]).
Let $a_{1}^{(N)},\ldots,a_{N}^{(N)}\in\mathbb{R}^{N}$ and $S_{N}$ be the symmetric group. We define the following symmetric measure on $\mathbb{R}^{N}$:
$\mu_{N}(\mathrm{d}\vec{x}):=\frac{1}{N!}\sum_{\sigma\in S_{N}}\prod_{j=1}^{N}\delta_{a_{j}^{(N)}}(\mathrm{d}x_{\sigma(j)}).$ (4.3)
Its one-dimensional marginal distribution is given by
@@
If there exists a compactly supported limiting measure $\rho$ and $\rho_{N}\to\rho$ in moments, we say that $\{\mu_{N}\}$ satisfies a Law of Large Numbers. Then one has
$\lim_{N\to\infty,\,\theta N\to c}G(\underbrace{u,0,\ldots,0}_{N\text{ entries}};\mu_{N})=\int_{\mathbb{R}}e^{ux}\rho^{(c)}(\mathrm{d}x)$ (4.5)
where $\rho^{(c)}$ and $\rho$ satisfies a Markov-Krein correspondence. We say that now $\{\mu_{N}\}$ is $c$-LLN–appropriate. In another words, $\{\mu_{N}\}$ satisfies a Law of Large Numbers implies that it is $c$-LLN–appropriate.
@@
The previous theorem is a special case of results in [3], while it is clear that it can also be deduced from our results (Theorem 1.1). In [3] they make the assumption that every measure encountered has compact support, and thus moments and cumulants can be defined without a problem, while we do not admit such an assumption. Also, in they paper they consider the Bessel generating function for more general measure than symmetric point masses, and give a limiting moment constraint for such measures to have limiting distributions. It is a future work to think about how one can relax such moment conditions to give a more general convergence constraint.
@@ 4.2. $c$-cumulants
Proposition 4.3 ([3] for the case $\rho$ is compactly supported).
@@
4.3. A positivity conjecture
In [3], it has been conjectured that for all $\theta>0$ and $N$-tuples $a_{1},\ldots,a_{N}$ and $b_{1},\ldots,b_{N}$, there exists a probability measure on $c_{1},\ldots,c_{N}$ such that
$\mathbb{E}\,\mathcal{B}_{\vec{c}}(\vec{x};N,\theta)=\mathcal{B}_{\vec{a}}(\vec{x};N,\theta)\mathcal{B}_{\vec{b}}(\vec{x};N,\theta)$ (4.9)
for any $\vec{x}\in\mathbb{C}$. This conjecture, as well as its generalisations, can be traced back to [51, 26, 49, 44]. Restricted to the case $\vec{x}=(x_{1},0,\ldots,0)$, this conjecture motivates a definition of a $\beta$-additive convolution, that is, there exists a measure being a $\beta$-additive convolution of two discrete measures
$\rho_{\vec{a}}=\frac{1}{N}\sum_{j=1}^{N}\delta_{a_{j}},\quad\rho_{\vec{b}}=\frac{1}{N}\sum_{j=1}^{N}\delta_{b_{j}},$ (4.10)
which is again a probability measure.
In the high temperature limit ($N\to\infty$ and $\beta N/2\to c$), there is a similar conjecture which has been mentioned in [45]. We state it in both the additive and multiplicative cases.
Conjecture 4.1.
For $c>0$ and any two probability measures $\rho_{1}$ and $\rho_{2}$ in $\mathcal{V}$ (c.f. (2.28)), there exists another probability measure $\mu\in\mathcal{V}$ such that
$\mu^{(c)}=\rho_{1}^{(c)}\ast\rho_{2}^{(c)},$ (4.11)
where $\ast$ denotes the regular additive convolution. Furthermore, if both $\rho_{1}$ and $\rho_{2}$ are defined on $\mathbb{R}^{+}$, there exists another probability measure $\nu\in\mathcal{V}$ (defined on $\mathbb{R}^{+}$) such that

### v3

@@ 3.4. Weak convergence in the high temperature regime
Theorem 2.9 tells us that $\delta$ can be arbitrary. We take $\delta=1$ and $c>1$. Summarising Lemma (3.9) and Lemma (3.12) we see that for any $\varepsilon\in(0,\frac{c-1}{5})$, $M_{\varepsilon}>0$ with $\rho([-M_{\varepsilon},M_{\varepsilon}])\in(\frac{1+5\varepsilon}{c},1]$ and $N>N_{0}$ one has
$\begin{split}&\left|\int_{\mathbb{R}}\rho_{N}^{(c_{N})}(x)e^{itx}-\int_{\mathbb{R}}\rho^{(c)}(x)e^{itx}\right|\leq\tilde{C}\left(\tilde{b}_{N}+\frac{|c_{N}-c|}{2e\varepsilon}\right)e^{\tilde{b}_{N}}+|c_{N}-c|B,\end{split}$ (3.99)
where
$\tilde{C}=\frac{\max\{\Gamma(c-\varepsilon),\Gamma(c+\varepsilon)\}|t|^{c+\varepsilon-1}}{2\pi}\cdot 2\sqrt{\frac{\pi}{\varepsilon}}e^{\frac{c}{2}\log(1+4M_{\varepsilon}^{2})+|t|},$ (3.100)
$\tilde{b}_{N}:=\frac{(c+\varepsilon)e^{\eta-1}}{\eta}\mathrm{d}_{W_{\eta}}(\rho_{N},\rho)+|c_{N}-c|\left(\frac{1}{2}\int_{\mathbb{R}}\log(1+s^{2})\rho(\mathrm{d}s)+4\right),$ (3.101)
$B=\left(\sup_{|w-c|\leq\epsilon}|\psi(w)|+|\log|t||+\frac{\pi}{2}\right)e^{\varepsilon\left(\sup_{|w-c|\leq\epsilon}|\psi(w)|+|\log|t||+\frac{\pi}{2}\right)}.$ (3.102)
The only $N$-dependent terms are $\tilde{b}_{N}$ and $|c_{N}-c|$, both of which are small when $N\to+\infty$. Thus we have shown the Fourier transforms of $\rho_{N}^{(c_{N})}$ and $\rho^{(c)}$ are close to each other, which by Lévy continuity implies the weak convergence. ∎
3.5. Proofs for Theorem 1.1 and 1.3
@@
and $\rho_{N}^{(c_{N})}$ is the distribution of the random mean of the Dirichlet process with parameter measure $c_{N}\rho$. By Section 3.3, in the limit where $c_{N}:=N\beta_{N}/2\to 0$ as $N\to+\infty$, the measure $\rho_{N}^{(c_{N})}$ converges weakly to $\rho$. Thus, by the dominated convergence theorem,
$\lim_{N\to+\infty}\int_{\mathbb{R}}e^{ux}\rho_{N}^{(c_{N})}(\mathrm{d}x)=\int_{\mathbb{R}}e^{ux}\rho(\mathrm{d}x),$ (3.105)
@@
Now for (1.24), its proof is almost identical, except that in the $N\to+\infty$ one has instead $c_{N}:=N\beta_{N}/2\to c>0$. Then by Section 3.4, one has the weak convergence $\rho_{N}^{(c_{N})}\to\rho^{(c)}$. Then applying the dominated convergence theorem gives us the result.
@@
3.6. Proof for Theorem 1.2
Theorem 1.2 is only a simple corollary of (3.99). Now since $a_{1}^{(N)},\ldots,a_{N}^{(N)}$ are random points, we have that $\rho^{(c_{N})}_{N}(\mathrm{d}x)$ is a random measure. Then extended $\eta$-Wasserstein distance $\mathrm{d}_{W_{\eta}}(\rho_{N},\rho)$ is then a random variable. For (3.99), both sides are random variables. One has
$\mathbb{E}\,\left|\int_{\mathbb{R}}e^{itx}(\rho_{N}^{(c_{N})}-\rho^{(c)})(\mathrm{d}x)\right|\leq\mathbb{E}\,\left[\tilde{C}\left(\tilde{b}_{N}+\frac{|c_{N}-c|}{2e\varepsilon}\right)e^{\tilde{b}_{N}}+|c_{N}-c|B\right].$ (3.108)
On the right hand side, the randomness only comes from $\mathrm{d}_{W_{\eta}}(\rho_{N},\rho)$ in $\tilde{b}_{N}$ which by (3.101) can be rewritten as
$\tilde{b}_{N}=\lambda\mathrm{d}_{W_{\eta}}(\rho_{N},\rho)+|c_{N}-c|\kappa$ (3.109)
and both $\lambda$ and $\kappa$ are deterministic. Therefore, with $|c_{N}-c|\to 0$, the convergence of the expectation on the right side of (3.108) boils down to the convergence of
$\mathbb{E}\,\left[\mathrm{d}_{W_{\eta}}(\rho_{N},\rho)\exp\mathrm{d}_{W_{\eta}}(\rho_{N},\rho)\right]\to 0,\qquad\sup_{n}\mathbb{E}\,\left[\exp\mathrm{d}_{W_{\eta}}(\rho_{N},\rho)\right]<+\infty,$ (3.110)
which is implied directly by Assumption 1.2.
@@ 4.1. Comparison with results in [7]
As mentioned in the introduction section, Benaych-Georges, Cuenca and Gorin [7] develop the notion of Bessel generating function $G_{N,\theta}(\vec{x};\mu)$ of a symmetric probability measure $\mu$ on $\mathbb{R}^{N}$, which is an integral transform with the multivariate Bessel function as the kernel:
$G_{N,\theta}(\vec{x};\mu):=\mathbb{E}\,\mathcal{B}_{\vec{a}}(\vec{u};\theta)=\int_{\mathbb{R}^{N}}\mathcal{B}_{\vec{a}}(\vec{u};\theta)\mu(\mathrm{d}\vec{a}).$ (4.1)
@@
Theorem 4.1 ( [7]).
Let $a_{1}^{(N)},\ldots,a_{N}^{(N)}\in\mathbb{R}$ be random and has a (uniformly in $n$) compactly supported empirical distribution
@@
If they are $c$-LLN–appropriate (for details check [7]) and $\rho_{N}\to\rho$ in moments, under suitable exponential tail control, one has
$\lim_{N\to+\infty,\,\theta N\to c}G(\underbrace{u,0,\ldots,0}_{N\text{ entries}};\mu_{N})=\int_{\mathbb{R}}e^{ux}\rho^{(c)}(\mathrm{d}x),\quad u\in\mathbb{C},$ (4.3)
where $\rho^{(c)}$ and $\rho$ satisfies a Markov-Krein correspondence.
@@
In the case $c>1$, the previous theorem is comparable to our results (Theorem 1.2), though there are several difference.
  1. (1)
    Theorem 4.1 makes the assumption that every measure encountered has compact support, and thus moments and cumulants can be defined without a problem, while we do not admit such an assumption.
  2. (2)
    Also, the convergence $\rho_{N}\to\rho$ together with the requirement of being $c$-LLN-appropriate are comparable to our Assumption 1.2.
  3. (3)
    Theorem 4.1 considers a general $u\in\mathbb{C}$, while we only have $u\in i\mathbb{R}$.
@@ 4.2. $c$-cumulants
Proposition 4.3 ([7] for the case $\rho$ being compactly supported).
@@
Acknowledgements
We thanks discussions with Tianshu Cong, Cesear Cuenca, Joon Lee, Pierre Mergny and Jiaming Xu and the proofreading of the first draft by Joon Lee and Pierre Mergny. JZ acknowledges financial support of FWO Odysseus grant No. G0DDD23N and FWO Fellowship Junior 1234325N.
Appendix A Proofs in Section 2
A.1. Proofs of Theorem 2.1 and Theorem 2.2
Although the arguments are already contained in [53], we provide them in line with our notions, only for the sake of completeness.
To do so, let us firstly consider the simple case when $\rho$ is a discrete measure, finitely supported on $\{x_{1},\ldots,x_{n}\}\subset\mathbb{R}$, given by
$\rho(\mathrm{d}x)=\sum_{j=1}^{n}a_{j}\delta_{x_{j}}(\mathrm{d}x)$ (A.1)
$\int_{\mathbb{R}^{n}}\frac{\mathrm{Dir}_{(ca_{1},\ldots,ca_{n})}(\mathrm{d}\sigma_{1},\ldots,\mathrm{d}\sigma_{n})}{(z-\sum_{j=1}^{n}x_{j}\sigma_{j})^{c}}$ $=\prod_{j=1}^{n}\frac{1}{(z-x_{j})^{a_{j}}}=\exp\left(-c\int_{\mathbb{R}}\log(z-s)\rho(\mathrm{d}s)\right).$ (A.2)
$(D_{c\rho}(\{x_{1}\}),\ldots,D_{c\rho}(\{x_{n}\}),D_{c\rho}(\mathbb{R}\setminus\{x_{1},\ldots,x_{n}\}))$ (A.3)
follows a Dirichlet distribution $\mathrm{Dir}_{(ca_{1},\ldots,ca_{n},0)}$. By convention of the Dirichlet distribution, this also implies its first $n$ entries follows a Dirichlet distribution $\mathrm{Dir}_{(ca_{1},\ldots,ca_{n})}$ while the last entry is $D_{c\rho}(\omega)(\mathbb{R}\setminus\{x_{1},\ldots,x_{n}\})=0$ almost surely. Hence
$\int_{\mathbb{R}^{n}}\varphi\left(\sum_{j=1}^{n}x_{j}\sigma_{j}\right)\mathrm{Dir}_{(ca_{1},\ldots,ca_{n})}(\mathrm{d}\sigma_{1},\ldots,\mathrm{d}\sigma_{n})$ $=\mathbb{E}\,\left[\varphi\left(\int_{\{x_{1},\ldots,x_{n}\}}xD_{c\rho}(\mathrm{d}x)\right)\right]$ (A.4)
holds for any bounded measurable function $\varphi$. Combining (A.2) and (A.4) we have
$\mathbb{E}\,\left[\left(z-\int_{\{x_{1},\ldots,x_{n}\}}xD_{c\rho}(\mathrm{d}x)\right)^{-c}\right]=\exp\left(-c\int_{\mathbb{R}}\log(z-s)\rho(\mathrm{d}s)\right),$ (A.5)
which holds for all $z\in\mathbb{C}\setminus\mathbb{R}$.
$\xi_{n}(\mathbb{R})=\left\{y_{j}:j\in J_{n}\right\},$ (A.6)
@@ Appendix A Proofs in Section 2
$\int_{\mathbb{R}}\varphi(x)\xi_{n,\sharp}\rho(\mathrm{d}x)=\int_{\mathbb{R}}\varphi(\xi_{n}(x))\rho(\mathrm{d}x),$ (A.8)
Lemma A.1 (Eq. (3.1) in [53]).
$\mathbb{E}\,\left[\left(z-\int_{\mathbb{R}}\varphi(\xi_{n}(x))D_{c\rho}(\mathrm{d}x)\right)^{-c}\right]=\exp\left(-c\int_{\mathbb{R}}\log(z-\varphi(s))\xi_{n,\sharp}\rho(\mathrm{d}s)\right).$ (A.9)
Proof.
$(\sigma_{j})_{j\in J_{n}}=(D_{c\rho}(\xi_{n}^{-1}(y_{j})))_{j\in J_{n}}$ (A.10)
$\int_{\mathbb{R}}\varphi(\xi_{n}(x))D_{c\rho}(\mathrm{d}x)=\sum_{j\in J_{n}}\varphi(y_{j})\sigma_{j}.$ (A.11)
The $j$-sum is finite and hence there is no convergence issue. Then by the Lauricella formula (A.2) we have
$\mathbb{E}\,\left[\left(z-\int_{\mathbb{R}}\varphi(\xi_{n}(x))D_{c\rho}(\mathrm{d}x)\right)^{-c}\right]$ $=\int_{\mathbb{R}^{n}}\frac{\mathrm{Dir}_{(ca_{j})_{j\in J_{n}}}(\mathrm{d}\vec{\sigma})}{(z-\sum_{j\in J_{n}}\varphi(y_{j})\sigma_{j})^{c}}=\prod_{j\in J_{n}}\frac{1}{(z-\varphi(y_{j}))^{ca_{j}}}.$ (A.12)
Since $a_{j}=\rho(\xi_{n}^{-1}(y_{j}))$, one has
$\begin{split}\prod_{j\in J_{n}}\frac{1}{(z-\varphi(y_{j}))^{ca_{j}}}&=\exp\left(-c\sum_{j\in J_{n}}\log(z-\varphi(y_{j}))\rho(\xi_{n}^{-1}(y_{j}))\right)\\ &=\exp\left(-c\int_{\mathbb{R}}\log(z-\varphi(\xi_{n}(s)))\rho(\mathrm{d}s)\right)=\exp\left(-c\int_{\mathbb{R}}\log(z-\varphi(s))\xi_{n,\sharp}\rho(\mathrm{d}s)\right).\end{split}$ (A.13)
Combining (A.12) and (A.13) finishes the proof. ∎
Now we are in a good shape to prove Theorem 2.1 and Theorem 2.2.
Proof of Theorem 2.1.
Using (A.9) and taking $z=1$ and $\varphi(x)=-|x|$, one obtains
$\mathbb{E}\,\left[\left(1+\int_{\mathbb{R}}|\xi_{n}(x)|D_{c\rho}(\mathrm{d}x)\right)^{-c}\right]=\exp\left(-c\int_{\mathbb{R}}\log(1+|s|)\xi_{n,\sharp}\rho(\mathrm{d}s)\right).$ (A.14)
$\mathbb{E}\,\left[\left(1+\lim_{n\to+\infty}\int_{\mathbb{R}}|\xi_{n}(x)|D_{c\rho}(\mathrm{d}x)\right)^{-c}\right]$ $=\exp\left(-c\int_{\mathbb{R}}\log(1+|s|)\rho(\mathrm{d}s)\right).$ (A.15)
On the other hand, assuming $\int_{\mathbb{R}}|x|D_{c\rho}(\mathrm{d}x)$ is finite almost surely, then by a similar argument the right side of (A.15) is non-zero, and thus we have (2.4). ∎
Proof of Theorem 2.2.
$\int_{\mathbb{R}}\frac{\rho^{(c)}(\mathrm{d}x)}{(z-ux)^{c}}=\exp\left(-c\int_{\mathbb{R}}\log(z-us)\rho(\mathrm{d}s)\right).$ (A.16)
For the existence of $\rho^{(c)}$ let us again start with (A.9) and take $\varphi(x)=ux$. Now one takes the limit $n\to+\infty$ on both sides, and again by the dominated convergence theorem we obtain
$\mathbb{E}\,\left[\left(z-u\lim_{n\to+\infty}\int_{\mathbb{R}}\xi_{n}(x)D_{c\rho}(\mathrm{d}x)\right)^{-c}\right]$ $=\exp\left(-c\int_{\mathbb{R}}\log(z-us)\rho(\mathrm{d}s)\right).$ (A.17)
A.2. Proof of Theorem 2.9
$\rho^{[M]}(\mathrm{d}s):=\rho((-\infty,-M])\delta_{-M}(\mathrm{d}s)+\mathbbm{1}_{(-M,M)}(s)\rho(\mathrm{d}s)+\rho([M,+\infty))\delta_{M}(\mathrm{d}s).$ (A.18)
The following Lemma has been used implicitly in [53, Section 7] without giving a proof. It is a special case of Corollary 2.4 (which is based on results in [21]), while here we provide a alternative elementary proof which does not rely on Corollary 2.4 and results in [21].
Lemma A.2.
Proof.
$\xi_{n}(x)=\begin{cases}\lfloor nx\rfloor/n,&x\in[0,n)\setminus\{M\},\\ M,&x=M,\\ n,&x>n,\\ -\xi_{n}(-x),&x<0,\end{cases}\qquad n>M.$ (A.19)
$\xi_{n}(\mathbb{R})=\{p/n:p=-n^{2},\ldots,-1,0,1,\ldots,n^{2}\}\cup\{\pm M\}.$ (A.20)
Let $\varphi$ be a continuous function with $|\varphi|\leq 1$. Then
$\mathbb{E}\,$ $\left[\varphi\left(\int\xi_{n}(x)D_{c\rho^{[M]}}(\mathrm{d}x)\right)\right]=\int\varphi\left(-M\sigma_{-}+\left(\sum_{\begin{subarray}{c}p\in\mathbb{Z}\cap(-nM,nM)\end{subarray}}\frac{p}{n}\sigma_{p}\right)+M\sigma_{+}\right){\mathrm{Dir}_{(c\vec{a})}(\mathrm{d}\vec{\sigma})}$ (A.21)
$a_{p}=\rho^{[M]}(\xi_{n}^{-1}(\{p/n\})),\quad p\in\mathbb{Z}\cap(-nM,nM),$ (A.22)
$a_{-}=\rho^{[M]}(\{-M\}),$ (A.23)
$a_{+}=\rho^{[M]}(\{M\}).$ (A.24)
Since $\xi_{n}^{-1}(\{p/n\})$ in contained in (M,M)(-M,M), we have
$a_{p}=\rho(\xi_{n}^{-1}(\{p/n\})),\quad p\in\mathbb{Z}\cap(-nM,nM),$ (A.25)
$a_{-}=\rho((-\infty,-M]),$ (A.26)
$a_{+}=\rho([M,+\infty)).$ (A.27)
Thus by the definition of the Dirichlet process, one has
$\mathbb{E}\,$ $\left[\varphi\left(\int\xi_{n}(x)P_{c\rho^{[M]}}(\mathrm{d}x)\right)\right]=\mathbb{E}\,\left[\varphi\left(\int\left(-M\mathbbm{1}_{x\leq-M}+\xi_{n}(x)\mathbbm{1}_{x\in(-M,M)}+M\mathbbm{1}_{x\geq-M}\right)D_{c\rho}(\mathrm{d}x)\right)\right].$ (A.28)
We then take the $N\to+\infty$ limit on both sides. On the left hand side, because of the continuity of $\varphi$ and the boundedness of $\xi_{n}(x)$ by $|x|$, one can push the limit inside both integrals and obtain
$\lim_{n\to+\infty}\mathbb{E}\,\left[\varphi\left(\int\xi_{n}(x)D_{c\rho^{[M]}}(\mathrm{d}x)\right)\right]=\mathbb{E}\,\left[\varphi\left(\int xD_{c\rho^{[M]}}(\mathrm{d}x)\right)\right].$ (A.29)
$\mathbb{E}\,\left[\varphi\left(\int xD_{c\rho^{[M]}}(\mathrm{d}x)\right)\right]=\mathbb{E}\,\left[\varphi\left(\int\left(-M\mathbbm{1}_{x\leq-M}+x\mathbbm{1}_{x\in(-M,M)}+M\mathbbm{1}_{x\geq-M}\right)D_{c\rho}(\mathrm{d}x)\right)\right].$ (A.30)
Taking $M\to+\infty$ limit on both sides gives
$\lim_{M\to\infty}\mathbb{E}\,\left[\varphi\left(\int xD_{c\rho^{[M]}(\mathrm{d}x)}\right)\right]=\mathbb{E}\,\left[\varphi\left(\int xD_{c\rho(\mathrm{d}x)}\right)\right].$ (A.31)
This proofs that $(\rho^{[M]})^{(c)}\to\rho^{(c)}$ weakly. ∎
Lemma A.3.
For all $c>0$, $u\in\mathbb{C}\setminus\{0\}$ and $x\in\mathbb{R}$, we have
$e^{ux}={u\Gamma(c)}\frac{1}{2\pi i}\int_{\mathrm{sgn}(u)\mathcal{C}}\mathrm{d}z\,\frac{e^{uz}}{(u(z-x))^{c}},$ (A.32)
where $\mathrm{sgn}$ is the sign function with complex arguments given by (2.27), and the contour $\mathcal{C}$ is the Hankel loop enclosing $(-\infty,x]$; see Fig. 1.
Proof.
$\frac{1}{\Gamma(c)}=\frac{1}{2\pi i}\int_{\mathcal{C}}\mathrm{d}z\,\frac{e^{z}}{z^{c}}.$ (A.33)
$\begin{split}\left|\int_{C_{R}}\frac{e^{uz}}{(u(z-x)^{c})}\right|&\leq\int_{\pi/2}^{\pi}\mathrm{d}\theta\frac{Re^{\mathrm{Re}\,(uRe^{i\theta})}}{|u(Re^{i\theta}-x)|^{c}}\\ &\leq\frac{\pi}{2}\frac{R}{|u(R-|x|)|^{c}}\sup_{\theta\in(\pi/2,\pi)}\exp(R(\mathrm{Re}\,u\cos\theta-\mathrm{Im}\,u\sin\theta)).\end{split}$ (A.34)
In the case $\mathrm{Re}\,u<0$, one can see that after the change of variable, the contour is deformed such that it starts and ends at the first (or fourth) quadrant if $\mathrm{Im}\,u\geq 0$ (or $\mathrm{Im}\,u<0$). A similar bound to (A.34) can be performed to see that applying a Cauchy theorem the contour can be deformed such that it encloses the branch point u(zx)u(z-x) with a branch cut going to $\infty$. We have
$\frac{1}{\Gamma(c)}=u\frac{1}{2\pi i}\int_{\mathcal{C}}\mathrm{d}z\,\frac{e^{u(z-x)}}{(u(z-x))^{c}}.$ (A.35)
A simplification gives us the result.
$\frac{1}{\Gamma(c)}=\frac{1}{2\pi i}\int_{-i\mathcal{C}}it\mathrm{d}z\,\frac{e^{it(z-x)}}{(it(z-x))^{c}}.$ (A.36)
The contour $i\mathcal{C}$ encloses the branch cut $i(-\infty,x]$.
Lemma A.4.
For $t\neq 0$ and $\varepsilon>0$ small enough,
$\int_{I_{t}}\exp(R(\varepsilon\cos\theta-t\sin\theta))\mathrm{d}\theta=O\left(\frac{e^{\varepsilon R}}{tR}\right),$ (A.37)
where $I_{t}=[\pi/2,\pi]$ if $t>0$, and $I_{t}=[\pi,3\pi/2]$ if $t<0$.
Proof.
$\begin{split}&\theta_{\max}=\arctan(-t/\varepsilon)\in\left(-\frac{\pi}{2},0\right),\quad\theta_{\min}=\arctan(-t/\varepsilon)+\pi\in\left(\frac{\pi}{2},\pi\right).\end{split}$ (A.38)
$\int_{a}^{b}\exp(R(\varepsilon\cos\theta-t\sin\theta))\mathrm{d}\theta=O\left(\frac{e^{R\phi(\pi)}}{R|\phi^{\prime}(\pi)|}\right).$ (A.39)
We also have
$\phi(\pi)=\varepsilon,\quad\phi^{\prime}(\pi)=-t\neq 0.$ (A.40)
Assembling the above gives the result.
With the previous two lemmas, one can now proof Theorem 2.9 fully rigorously.
Proof of Theorem 2.9.
Step I: Compactly supported $\rho$. By Theorem 2.6, the left side of (2.30) is finite. By Lemma A.3 we have
$\int_{\mathbb{R}}\rho^{(c)}(\mathrm{d}x)e^{ux}=\int_{-M}^{M}\rho^{(c)}(\mathrm{d}x)\,{u\Gamma(c)}\frac{1}{2\pi i}\int_{\mathrm{sgn}(u)\mathcal{C}}\mathrm{d}z\frac{e^{uz}}{(u(z-x))^{c}}.$ (A.41)
$\begin{split}\int_{\mathbb{R}}\rho^{(c)}(\mathrm{d}x)e^{ux}&=y\Gamma(c)\frac{1}{2\pi i}\int_{\mathrm{sgn}(u)\mathcal{C}}\mathrm{d}z\,e^{uz}\int_{-M}^{M}\frac{\rho^{(c)}(\mathrm{d}x)}{(u(z-x))^{c}}\\ &=u\Gamma(c)\frac{1}{2\pi i}\int_{\mathrm{sgn}(u)\mathcal{C}}\mathrm{d}z\,\exp\left(uz-c\int_{\mathbb{R}}\log(u(z-s))\rho(\mathrm{d}s)\right),\end{split}$ (A.42)
$\log(u(z-s))=\log u+\log(z-s)+2m\pi i$ (A.43)
$\int_{\mathbb{R}}\rho^{(c)}(\mathrm{d}x)e^{ux}$ $=\frac{\Gamma(c)}{u^{c-1}}\frac{1}{2\pi i}\int_{\mathrm{sgn}(u)\mathcal{C}}\mathrm{d}z\,\exp\left(uz-c\int_{\mathbb{R}}\log(z-s)\rho(\mathrm{d}s)\right).$ (A.44)
Step II: $u=it$ with $t>0$. In the special case where $\rho$ is compactly supported and $u=it$ with $t>0$, we need to deform the contour $\mathrm{sgn}(u)\mathcal{C}$ to $\mathbb{R}-i\delta$, for every $\delta>0$. For a graphical illustration of the deformation, see Fig. 2.